OpenStudy (anonymous):

Help Please *MEDAL and FAN!* What is the y-intercept of the line represented by the equation? y = 7x + 2

OpenStudy (anonymous):

@Luigi0210 @undeadknight26

OpenStudy (luigi0210):

Y-intercept is when x=0.

OpenStudy (ribhu):

an eqn of the form y=mx+c intercept is c directly

OpenStudy (ribhu):

or else put x=0, intercept would be on y axis,i.e the value of y at x=0

OpenStudy (anonymous):

ok

OpenStudy (ribhu):

here in this case compare the equation y=mx+c

OpenStudy (anonymous):

can some one help me with my question

OpenStudy (luigi0210):

Y-intercept, in short is xD \(\Large y=7(0)+2 \) :P

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'd give the medal to both ... but idk who so my answer is 0?

OpenStudy (ribhu):

know the intercept is 2, the value of y at which x is zero

OpenStudy (anonymous):

mhm

OpenStudy (ribhu):

@FateRocks8D u get it!!

OpenStudy (anonymous):

o.o !!!! yesh

OpenStudy (ribhu):

where are our medals? and fans? lolz

OpenStudy (anonymous):

*gives medals to both but I cant only give for one :S ... and I fanned -.- mr.pushy (jk jk jk jk

OpenStudy (anonymous):

don't kill me o.o @Luigi0210

OpenStudy (luigi0210):

xD I wouldn't \(\Large \color{red}{\heartsuit^{\heartsuit^{\color{purple}{\heartsuit}}}} \) ;P

OpenStudy (anonymous):

Awe <3